Circular No. 7561

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 A1 (LINEAR) M. Blythe reports the discovery by LINEAR of a new comet (discovery

observation below). Following posting on the NEO Confirmation Page, several other CCD observers confirmed the object’s cometary nature: M. Dawson (Luxembourg) found the object to be diffuse with a 9′′ coma on Jan. 13.0 UT; Jan. 14.2 images taken by L. Kornoš and P. Kolény (Modra) and by L. Šarounová (Ondřejov) showed coma diameters of ∼ 15′′ (m_{1} = 16.3) and ∼ 20′′ (m_{1} = 16.4), respectively; and images taken on Jan. 14.9 by J. Tichá and M. Tichý (Kleť) showed a diffuse coma of diameter 17′′ and a faint 30′′ tail in p.a. 200^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 7.46627  12^{h}01^{m}58.^{s}99  +42^{o}02′17′′.7  17.7

Full astrometric data and the following preliminary parabolic orbital elements are given on MPEC 2001-A54:

T = 2000 Sept.25.951 TT  ω = 112.^{o}926
Ω = 338.773 }2000.0
q = 2.52385 AU  i = 61.786
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 11 11^{h}57.^{m}71 +41^{o}56.^{′}5 2.132 2.773 121.^{o}3 17.^{o}6 16.6
21 11 42.77 +41 35.3 2.074 2.819 130.9 15.3 16.6
31 11 24.12 +40 57.6 2.039 2.867 140.2 12.7 16.6
Feb. 10 11 03.10 +39 52.7 2.034 2.918 148.0 10.3 16.7
20 10 41.60 +38 15.4 2.062 2.972 152.0  9.0 16.8
Mar. 2 10 21.59 +36 07.5 2.126 3.027 150.3  9.3 16.9
12 10 04.60 +33 37.3 2.226 3.086 144.0 10.9 17.1

V445 PUPPIS W. Liller, Viña del Mar, Chile, writes: “Low-resolution CCD spectra

taken on Jan. 6.09, 10.10, and 14.09 UT show a slight weakening of the Fe II lines recorded by me on Jan. 3.23 (IAUC 7556). Otherwise, the four spectra are all very similar. While the lack of hydrogen emission is unusual, other cataclysmic stars such as CR Boo, CP Eri, and V803 Cen are similarly hydrogen-deficient. However, unlike V445 Pup, these stars, are all hot, blue objects showing He lines in emission and never Ca I or Na I.”

2001 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green